# On the Bergman metric of symmetric spaces

### v1

@@ Abstract.
The authors are supported by INdAM and GNSAGA - Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni, by GOACT and MAPS- Funded by Fondazione di Sardegna, by the Italian Ministry of Education, University and Research through the PRIN 2022 project "Developing Kleene Logics and their Applications" (DeKLA), project code: 2022SM4XC8. The second author acknowledges the support and hospitality of Cergy University, where part of this work was conducted during a research stay.
@@ 1. Introduction
In particular, $g_{\mathbb{B}^{n}}$ has constant holomorphic sectional curvature $\frac{4}{n+1}$, as it differs only by a factor from the hyperbolic metric on $\mathbb{B}^{n}$. The interested reader can refer to [11] for a detailed exposition on this topic.
The studies accomplished so far highlight how $g_{\Omega}$ encodes the intrinsic geometric behavior of $\Omega$. One key example of such results has been given by Q.-K. Lu in 1965 (see [15]), where he establishes that if $g_{\Omega}$ is complete with constant holomorphic sectional curvature then $\Omega$ is biholomorphic to $\mathbb{B}^{n}$. In 2025, X. Huang, S.-Y. Li and J. N. Treuer (see [7]) proved that if $g_{\Omega}$ has constant holomorphic sectional curvature then it must be negative, and assuming pseudoconvexity instead of completeness, proved that $\Omega$ is biholomorphic to $\mathbb{B}^{n}$ with possibly a pluripolar set removed. Observe that this uniformization generalizes Lu’s one, as in 1955 Bremermann proved (see [2]) that if $g_{\Omega}$ is complete then $\Omega$ is pseudoconvex, while the converse is not true: the classical counterexample being the pseudoconvex domain obtained by removing a complex hyperplane from $\mathbb{B}^{n}$. Starting from [7], in 2025 P. Ebenfelt, J. N. Treuer and M. Xiao (see [5]) maintained only the curvature hypothesis and obtained that $\Omega$ is biholomorphic to $\mathbb{B}^{n}$ with possibly a set of zero Lebesgue measure removed, over which $A^{2}$ functions extend holomorphically to $\mathbb{B}^{n}$.
@@
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 1$, be a bounded domain. If $\nabla R^{\Omega}=0$ and $\Omega$ is pseudoconvex, then there exists $\widetilde{\Omega}\subset\mathbb{C}^{n}$ bounded symmetric domain, ad $E\subset\widetilde{\Omega}$ pluripolar and closed in $\widetilde{\Omega}$ such that $\Omega\cong\widetilde{\Omega}\smallsetminus E$.
@@ 2. Proofs of the main results
In the proofs of Theorem 1.1 and 1.2, we need the following technical results. A striking difference is that, for the ball $\mathbb{B}^{n}$, positive scalings of the Bergman metric are projectively induced, whereas for a general bounded symmetric domain, projective induced-ness holds only for scalings governed by the Wallach set; cf. Proposition 2.2(A).
@@
It is well-known (see, e.g., [14] Sec. 3.4) that $\mathcal{B}^{\mathcal{S}}$ is a full Kähler immersion.
Step 1: $\mathcal{B}^{\mathcal{S}}$ is injective. Assume $\mathcal{B}^{\mathcal{S}}(p)=\mathcal{B}^{\mathcal{S}}(q)$ for some $p,q\in\Omega$. Then there exists $\lambda\in\mathbb{C}^{*}$ such that $\phi_{j}(p)=\lambda\,\phi_{j}(q)$ for all $j\in\mathbb{N}^{*}$. Let $K_{\Omega}$ be the Bergman kernel of $\Omega$. Using the expansion (1)
@@
and $D_{FS}([Z],[W])=0$ if and only if $[Z]=[W]$ (by equality in Cauchy–Schwarz). Moreover, by Calabi’s hereditary property of the diastasis (see, e.g., [3]), the diastasis of $(\Omega,cg_{\Omega})$ is given by the pullback of $D_{FS}$ via $F$, i.e.
@@
whenever the diastasis is defined. Assume $F(p)=F(q)$. Then $D_{FS}(F(p),F(q))=0$, hence $D^{c\Omega}(p,q)=0$. Since the diastasis scales linearly with the metric, i.e. $D^{c\Omega}=c\,D^{\Omega}$, we also have $D^{\Omega}(p,q)=0$. Again by Calabi’s hereditary property of the diastasis (see, e.g., [3]), if $\mathcal{B}^{\mathcal{S}}$ is a Bergman–Bochner map as in (4), then
@@
  • $\gamma_{k}$ the genus of $\Omega_{k}$.
@@
$\Gamma_{\Omega}:=\max_{1\leq k\leq s}\left\{\frac{\inf W_{c}(\Omega_{k})}{\gamma_{k}}\right\}\in\mathbb{R}_{\geq 0}.$
If $\delta>\Gamma_{\Omega}$, then $\delta>\inf W_{c}(\Omega_{k})/\gamma_{k}$ for every $k$, hence $\delta\,\gamma_{k}\in W_{c}(\Omega_{k})$ for all $k$. By [13, Theorem 2] it follows that, for every $k=1,\dots,s$, the scaled Bergman manifold $(\Omega_{k},\delta g_{\Omega_{k}})$ admits a full Kähler immersion
@@
(B). Fix $k\in\{1,\dots,s\}$ and let $N_{\Omega_{k}}$ be the generic norm of $\Omega_{k}$. It is well-known (see, e.g., [18, Sec. 1.3]) that on the diagonal one has
@@
and there are $m_{1},\dots,m_{r}$ polynomials of of bidegrees $(1,1),\dots,(r,r)$ on $\Omega_{k}\times\text{conj}(\Omega_{k})$ such that
$N_{\Omega_{k}}(z,\bar{z})=1+\sum_{\ell=1}^{r}(-1)^{\ell}\,m_{\ell}(z,\bar{z}).$ (7)
@@
cannot contain non-constant purely holomorphic or antiholomorphic terms. Applying this with $\alpha=-\gamma_{k}$ and using (6)–(7), we conclude that the real analytic expansion of $z\mapsto K_{\Omega_{k}}(z,z)$ at 0 contains no non-constant purely holomorphic nor purely antiholomorphic terms. Finally, since the Bergman kernel factorizes on products, writing $z=(z^{(1)},\dots,z^{(s)})\in\Omega_{1}\times\cdots\times\Omega_{s}$
$K_{\Omega}(z,z)=\prod_{k=1}^{s}K_{\Omega_{k}}\bigl(z^{(k)},z^{(k)}\bigr),$
hence the same conclusion holds for $K_{\Omega}(z,z)$. This proves (B). ∎
@@
Let $\pi\colon(X,\pi^{*}g_{\Omega})\to(\Omega,g_{\Omega})$ be the universal covering. Since $g_{\Omega}$ is complete and $\nabla R^{\Omega}=0$, $(X,\pi^{*}g_{\Omega})$ is simply connected, complete and locally symmetric, hence a (globally) Hermitian symmetric space of noncompact type. Therefore there exist a bounded symmetric domain $\widetilde{\Omega}\subset\mathbb{C}^{n}$ and a constant $\lambda>0$ together with a biholomorphic isometry
@@
Let $q\in\Omega$ and $\eta_{q}\colon[0,1]\to\Omega$ be a simple continuous path with $\eta_{q}(0)=p,\ \eta_{q}(1)=q$. Repeatedly applying Calabi’s rigidity theorem [3, Theorem 9], we find $A_{q}$ connected neighborhood of $\eta_{q}([0,1])\cup U$ and a Kähler immersion
@@
Furthermore, if $r\in\Omega\smallsetminus\{q\}$ and $A_{q}\cap A_{r}\neq\emptyset$, we have
@@
$\widetilde{f}\colon(\Omega,kg_{\Omega})\to(\widetilde{\Omega},k\lambda g_{\widetilde{\Omega}}):\widetilde{f}(s):=\widetilde{f}_{s}(s)$
@@
In particular, $\widetilde{f}$ is injective because $T\circ\mathcal{B}^{\Omega}$ is injective, by Proposition 2.1. By the open mapping theorem and the pseudoconvexity of $\Omega$, $W:=\widetilde{f}(\Omega)$ is a pseudoconvex domain in $\widetilde{\Omega}$ and $\widetilde{f}\colon(\Omega,g_{\Omega})\to(W,\lambda g_{\widetilde{\Omega}}|_{W})$ is a biholomorphic isometry. By the biholomorphic invariance of the Bergman metric (see, e.g., [11, Proposition 1.1.18]), we also get
@@
$\partial\bar{\partial}\left(\log(K_{W}(z,z))-\log(|\phi_{1}(z)|^{2})-\lambda\log(K_{\widetilde{\Omega}}(z,z))\right)=0\ ,\quad z\in W$ (9)
that is, $z\mapsto\log(K_{W}(z,z))-\log(|\phi_{1}|^{2}(z))-\lambda\log(K_{\widetilde{\Omega}}(z,z))$ is pluriharmonic on $W$. Consequently there exist $\Lambda\subseteq W$ neighborhood of 0, $h\in\mathcal{O}(\Lambda),\ h=\sum_{|\alpha|\geq 0}a_{\alpha}z^{\alpha}$, such that for all $z\in\Lambda$, $\phi_{1}(z)\neq 0$ and
@@
$-\log(K_{\widetilde{\Omega}}(z,z))=\log\left(\frac{1}{\text{vol}(\widetilde{\Omega})}\right)\sum_{k=1}^{\infty}\frac{\Big(1-\text{vol}(\widetilde{\Omega})K_{\widetilde{\Omega}}(z,z)\Big)^{k}}{k}.$
We infer by Proposition 2.2 that the real analytic expansion of $z\mapsto-\lambda\log(K_{\widetilde{\Omega}}(z,z))$ at 0 does not contain non-constant purely holomorphic or antiholomorphic terms. Secondly, by the properties of the Bergman kernel
@@
$L_{1}:=\partial W\cap Z(h_{0})\ ,\quad L_{2}:=\partial W\cap(\widetilde{\Omega}\smallsetminus Z(h_{0}))$
so that $E=L_{1}\cup L_{2}$, and since $L_{1}$ is pluripolar, it only suffices to prove that $L_{2}$ is pluripolar. Let $z_{0}\in L_{2}$ and $B_{\varepsilon}(z_{0})\subset\widetilde{\Omega}\smallsetminus Z(h_{0})$. Then $\frac{1}{h_{0}}\in\mathcal{O}(B_{\varepsilon}(z_{0})\cap W)$, so there exist $\varepsilon^{\prime}\in(0,\varepsilon)$ and $t\in\mathbb{R}^{>0}$ such that
@@
Thus, if $U_{z_{0}}$ is a neighborhood of $z_{0}$, by [17, Lemma 11]: $P_{z_{0}}:=U_{z_{0}}\smallsetminus W$ is pluripolar. In particular, we now show
@@
$\partial W\cap U_{z_{0}}\subseteq U_{z_{0}}\cap(\mathbb{C}^{n}\smallsetminus W)=P_{z_{0}}.$
On the other hand, if by contradiction there exist $a\in P_{z_{0}}\cap(\mathbb{C}^{n}\smallsetminus\overline{W})$, then there exists $U_{a}$ neighborhood of $a$ such that
$U_{a}\subset P_{z_{0}}\cap(\mathbb{C}^{n}\smallsetminus\overline{W})\subset\partial W\cap U_{z_{0}}\subseteq U_{z_{0}}\cap(\mathbb{C}^{n}\smallsetminus W)=P_{z_{0}}$
which is absurd, because it is well-known (see, e.g., [4, Theorem 4.17]) that a pluripolar set has zero Lebesgue measure. It is then proved that $L_{2}$ is locally pluripolar, hence pluripolar by [9, Josefson’s theorem], as desired. To conclude the proof, we now show $\widetilde{\Omega}\smallsetminus E=W$. Clearly $\widetilde{\Omega}\smallsetminus E\supseteq W$. Conversely, just notice that
$\widetilde{\Omega}\smallsetminus E\subset W\cup(\mathbb{C}^{n}\smallsetminus\overline{W})$
and if $(\widetilde{\Omega}\smallsetminus E)\cap(\mathbb{C}^{n}\smallsetminus\overline{W})\neq\emptyset$, we would have a separation of $\widetilde{\Omega}\smallsetminus E$, which contradicts the well-known fact (see, e.g., [4, Corollary 5.26]) that a pluripolar and relatively closed subset of a domain does not disconnect the domain.

### v2

@@ Abstract.
The authors are supported by INdAM and GNSAGA - Gruppo Nazionale per le Strutture Algebriche, Geometriche e le loro Applicazioni and by the project ProBiKi of Fondazione di Sardegna (Italy). The second author acknowledges the support and hospitality of Cergy University, where part of this work was conducted during a research stay.
@@ 1. Introduction
In particular, $g_{\mathbb{B}^{n}}$ has constant holomorphic sectional curvature $\frac{4}{n+1}$, as it differs only by the factor $n+1$ from the hyperbolic metric on $\mathbb{B}^{n}$. The interested reader can refer to [12] for a detailed exposition on this topic.
The studies accomplished so far highlight how $g_{\Omega}$ encodes the intrinsic geometric behavior of $\Omega$. One key example of such results has been given by Q.-K. Lu in 1966 (see [16]), where he establishes that if $g_{\Omega}$ is complete with constant holomorphic sectional curvature then $\Omega$ is biholomorphic to $\mathbb{B}^{n}$. In 2025, X. Huang, S.-Y. Li and J. N. Treuer (see [7]) proved that if $g_{\Omega}$ has constant holomorphic sectional curvature then it must be negative, and assuming pseudoconvexity instead of completeness, proved that $\Omega$ is biholomorphic to $\mathbb{B}^{n}$ with possibly a pluripolar set removed. Observe that this uniformization generalizes Lu’s one, as in 1955 Bremermann proved (see [2]) that if $g_{\Omega}$ is complete then $\Omega$ is pseudoconvex, while the converse is not true: the classical counterexample being the pseudoconvex domain obtained by removing a complex hyperplane from $\mathbb{B}^{n}$. Starting from [7], in 2025 P. Ebenfelt, J. N. Treuer and M. Xiao (see [5]) maintained only the curvature hypothesis and obtained that $\Omega$ is biholomorphic to $\mathbb{B}^{n}$ with possibly a set of zero Lebesgue measure removed, over which $A^{2}$ functions extend holomorphically to $\mathbb{B}^{n}$.
@@
Let $\Omega\subset\mathbb{C}^{n}$, $n\geq 1$, be a bounded domain. If $\nabla R^{\Omega}=0$ and $\Omega$ is pseudoconvex, then there exists $\widetilde{\Omega}\subset\mathbb{C}^{n}$ bounded symmetric domain, and $E\subset\widetilde{\Omega}$ pluripolar and closed in $\widetilde{\Omega}$ such that $\Omega\cong\widetilde{\Omega}\setminus E$.
@@ 2. Proofs of the main results
In the proofs of Theorem 1.1 and 1.2, we need the following technical results about the Bergman metric of bounded domains.
We start with a general fact, valid for any bounded domain $\Omega\subset\mathbb{C}^{n}$, concerning the injectivity of Kähler immersions of positive scalings of the Bergman metric into $\mathbb{CP}^{\infty}$.
@@
It is well-known (see, e.g., [15, Sec. 3.4]) that $\mathcal{B}^{\mathcal{S}}$ is a full Kähler immersion.
Step 1: $\mathcal{B}^{\mathcal{S}}$ is injective. Assume by contradiction $\mathcal{B}^{\mathcal{S}}(p)=\mathcal{B}^{\mathcal{S}}(q)$ for some $\in\Omega, pqp\neq q$, so there exists $\lambda\in\mathbb{C}^{*}$ such that $\phi_{j}(p)=\lambda\,\phi_{j}(q)$ for all $j\in\mathbb{N}^{*}$. Let $K_{\Omega}$ be the Bergman kernel of $\Omega$. Using the expansion (1)
@@
and $D_{FS}([Z],[W])=0$ if and only if $[Z]=[W]$ (by equality in Cauchy–Schwarz). Moreover, by Calabi’s hereditary property of the diastasis (see, e.g., [3]), the diastasis of $c\Omega:=(\Omega,cg_{\Omega})$ is given by the pullback of $D_{FS}$ via $F$, i.e.
@@
whenever the diastasis is defined. Assume $F(p)=F(q)$. Then
$D_{FS}(F(p),F(q))=D^{c\Omega}(p,q)=0.$
Since the diastasis scales linearly with the metric, i.e. $D^{c\Omega}=c\,D^{\Omega}$, we also have $D^{\Omega}(p,q)=0$. Again by Calabi’s hereditary property of the diastasis (see, e.g., [3]), if $\mathcal{B}^{\mathcal{S}}$ is a Bergman–Bochner map as in (4), then
@@
In the next proposition we collect three properties that will be repeatedly used in the sequel. The striking point in (A) is that, contrary to the ball $\mathbb{B}^{n}$ for which every positive multiple of the Bergman metric is projectively induced, for a general bounded symmetric domain projective inducedness occurs only for scalings prescribed by the Wallach set. On the other hand, (B) and (C) reflect features that are shared with $\mathbb{B}^{n}$.
@@
  • (C)
    For any $z\in\Omega$ there is a constant $C_{z}\in\mathbb{R}^{>0}$ dependent on $z$ such that $|K_{\Omega}(z,w)|\leq C_{z},\,\forall\,w\in\Omega$.
@@
  • $\gamma_{k}$ the genus of $\Omega_{k}$;
  • $r_{k}$ the rank of $\Omega_{k}$.
@@
$\Gamma_{\Omega}:=\max_{1\leq k\leq s}\left\{\frac{\inf W_{c}(\Omega_{k})}{\gamma_{k}}\right\}\in\mathbb{R}_{\geq 0}.$ (6)
If $\delta>\Gamma_{\Omega}$, then $\delta>\inf W_{c}(\Omega_{k})/\gamma_{k}$ for every $k$, hence $\delta\,\gamma_{k}\in W_{c}(\Omega_{k})$ for all $k$. By [14, Theorem 2] it follows that, for every $k=1,\dots,s$, the scaled Bergman manifold $(\Omega_{k},\delta g_{\Omega_{k}})$ admits a full Kähler immersion
@@
(B) Fix $k\in\{1,\dots,s\}$ and let $N_{\Omega_{k}}$ be the generic norm of $\Omega_{k}$. It is well-known (see, e.g., [19, Sec. 1.3]) that on the diagonal one has
@@
and there are $m_{1},\dots,m_{r_{k}}$ polynomials of of bidegrees $(1,1),\dots,(r_{k},r_{k})$ on $\Omega_{k}\times\text{conj}(\Omega_{k})$ respectively, such that
$N_{\Omega_{k}}(z,\bar{z})=1+\sum_{\ell=1}^{r_{k}}(-1)^{\ell}\,m_{\ell}(z,\bar{z}).$ (8)
@@
cannot contain non-constant purely holomorphic or antiholomorphic terms. Applying this with $H=H_{k},\,\alpha=-\gamma_{k}$, and using (7)–(8) we conclude that the real analytic expansion of $z\mapsto K_{\Omega_{k}}(z,z)$ at 0 contains no non-constant purely holomorphic nor purely antiholomorphic terms. Finally, recall that the Bergman kernel factorizes on products, i.e. writing $z=(z^{(1)},\dots,z^{(s)})\in\Omega_{1}\times\cdots\times\Omega_{s}$
$K_{\Omega}(z,w)=\prod_{k=1}^{s}K_{\Omega_{k}}\bigl(z^{(k)},w^{(k)}\bigr).$ (9)
Hence the same conclusion holds for $K_{\Omega}(z,z)$. This proves (B).
(C) Assume $\Omega$ is irreducible. Fix $z\in\Omega$. Since $\Omega$ is symmetric, $\operatorname{Aut}(\Omega)$ acts transitively; hence there exists $\varphi^{z}\in\operatorname{Aut}(\Omega)$ such that $\varphi^{z}(z)=0$. Let $J_{\varphi^{z}}(\xi):=\det_{\mathbb{C}}(d(\varphi^{z})_{\xi})$ be the complex Jacobian determinant of $\varphi^{z}$. In its Harish–Chandra realization, $\Omega$ is a bounded (positive) circular domain with $0\in\Omega$, hence by [10, Theorem 1] every $\varphi\in\operatorname{Aut}(\Omega)$ extends holomorphically to $\overline{\Omega}$; in particular, $\varphi^{z}$ (and thus $J_{\varphi^{z}}$) extends holomorphically to $\overline{\Omega}$. Therefore, being $\overline{\Omega}$ compact and $\varphi^{z}$ biholomorphic
$M_{z}:=\sup_{\xi\in\Omega}|J_{\varphi^{z}}(\xi)|\in\mathbb{R}^{>0}$
We now use the transformation rule for the Bergman kernel under biholomorphisms: for every $\psi\in\operatorname{Aut}(\Omega)$ and all $u,v\in\Omega$,
$K_{\Omega}(u,v)=J_{\psi}(u)\,\overline{J_{\psi}(v)}\,K_{\Omega}(\psi(u),\psi(v)).$ (10)
Applying (10) with $\psi=\varphi^{z}$ and $u=z$, we obtain
$K_{\Omega}(z,w)=J_{\varphi^{z}}(z)\,\overline{J_{\varphi^{z}}(w)}\,K_{\Omega}(0,\varphi^{z}(w)),\qquad w\in\Omega.$ (11)
Next we claim that $K_{\Omega}(0,\cdot)$ is constant. Indeed, since in its Harish–Chandra realization $\Omega$ is circled, for every $\theta\in\mathbb{R}$ the rotation $r_{\theta}(w)=e^{\sqrt{-1}\theta}w$ belongs to $\operatorname{Aut}(\Omega)$ and satisfies $J_{r_{\theta}}\equiv e^{\sqrt{-1}n\theta}$. Using (10) with $\psi=r_{\theta}$ and $u=0$ gives
$K_{\Omega}(0,w)=K_{\Omega}(0,e^{\sqrt{-1}\theta}w),\qquad\forall\,\theta\in\mathbb{R},\ \forall\,w\in\Omega.$
Since $w\mapsto K_{\Omega}(0,w)$ is antiholomorphic, its power series at 0 has the form $K_{\Omega}(0,w)=\sum_{\alpha}c_{\alpha}\,\overline{w^{\alpha}}$; the above invariance forces $c_{\alpha}=0$ for all $|\alpha|\geq 1$. Thus, $K_{\Omega}(0,w)\equiv K_{\Omega}(0,0)=\frac{1}{{\rm vol}(\Omega)}$. Plugging this into (11) yields
$|K_{\Omega}(z,w)|=\frac{|J_{\varphi^{z}}(z)|}{{\rm vol}(\Omega)}\,|J_{\varphi^{z}}(w)|\leq\frac{|J_{\varphi^{z}}(z)|}{{\rm vol}(\Omega)}\,M_{z},\qquad\forall\,w\in\Omega.$
Setting $C_{z}:=\dfrac{|J_{\varphi^{z}}(z)|}{{\rm vol}(\Omega)}\,M_{z}\in\mathbb{R}^{>0}$ proves (C) in the irreducible case, hence in general by (9). ∎
@@
Let $\pi\colon(X,\pi^{*}g_{\Omega})\to(\Omega,g_{\Omega})$ be the universal covering. Since $g_{\Omega}$ is complete and $\nabla R^{\Omega}=0$, $(X,\pi^{*}g_{\Omega})$ is simply connected, complete and locally symmetric, hence a (globally) Hermitian symmetric space of noncompact type. Therefore there exist a bounded symmetric domain $\widetilde{\Omega}\subset\mathbb{C}^{n}$ and a constant $\lambda\in\mathbb{R}^{>0}$ together with a biholomorphic isometry
@@
Let $q\in\Omega$ and $\eta_{q}\colon[0,1]\to\Omega$ be a simple continuous path with $\eta_{q}(0)=p,\ \eta_{q}(1)=q$. Repeatedly applying Calabi’s rigidity theorem, we find $A_{q}$ connected neighborhood of $\eta_{q}([0,1])\cup U$ and a Kähler immersion
@@
Furthermore, if $r\in\Omega\setminus\{q\}$ and $A_{q}\cap A_{r}\neq\emptyset$, we have
@@
$\widetilde{f}\colon(\Omega,kg_{\Omega})\to(\widetilde{\Omega},k\lambda g_{\widetilde{\Omega}}),\ \widetilde{f}(s):=\widetilde{f}_{s}(s)$
@@
In particular, $\widetilde{f}$ is injective because $T\circ\mathcal{B}^{\Omega}$ is injective, by Proposition 2.1. By the open mapping theorem and the pseudoconvexity of $\Omega$, $W:=\widetilde{f}(\Omega)$ is a pseudoconvex domain in $\widetilde{\Omega}$ and $\widetilde{f}\colon(\Omega,g_{\Omega})\to(W,\lambda g_{\widetilde{\Omega}}|_{W})$ is a biholomorphic isometry. By the biholomorphic invariance of the Bergman metric, we also get
@@
$\partial\bar{\partial}\left(\log(K_{W}(z,z))-\log(|\phi_{1}(z)|^{2})-\lambda\log(K_{\widetilde{\Omega}}(z,z))\right)=0\ ,\quad z\in\Lambda^{\prime}$ (13)
that is, $z\mapsto\log(K_{W}(z,z))-\log(|\phi_{1}|^{2}(z))-\lambda\log(K_{\widetilde{\Omega}}(z,z))$ is pluriharmonic on $\Lambda^{\prime}$. Consequently there exist $\Lambda\subseteq\Lambda^{\prime}$ neighborhood of 0, $h\in\mathcal{O}(\Lambda),\ h=\sum_{|\alpha|\geq 0}a_{\alpha}z^{\alpha}$, such that
@@
$-\log(K_{\widetilde{\Omega}}(z,z))=\log\big({\text{vol}(\widetilde{\Omega})}\big)+\sum_{k=1}^{\infty}\frac{\Big(1-\text{vol}(\widetilde{\Omega})K_{\widetilde{\Omega}}(z,z)\Big)^{k}}{k}.$
We infer by Proposition 2.2 (B) that the real analytic expansion of $z\mapsto-\lambda\log(K_{\widetilde{\Omega}}(z,z))$ at 0 does not contain non-constant purely holomorphic or antiholomorphic terms. Secondly, by the properties of the Bergman kernel
@@
$E\cap U\subset\{z\in U\ |\ u(z)=-\infty\}.$
Set
$L_{1}:=\partial W\cap Z(h_{0})\ ,\quad L_{2}:=\partial W\cap(\widetilde{\Omega}\setminus Z(h_{0}))$
so that $E=L_{1}\cup L_{2}$, and since $L_{1}$ is pluripolar, it only suffices to prove that $L_{2}$ is pluripolar. Let $z_{0}\in L_{2}$ and $B_{\varepsilon}(z_{0})\subset\widetilde{\Omega}\setminus Z(h_{0})$. Then $\frac{1}{h_{0}}\in\mathcal{O}(B_{\varepsilon}(z_{0})\cap W)$, so there exist $\varepsilon^{\prime}\in(0,\varepsilon)$ and $t\in\mathbb{R}^{>0}$ such that
@@
Thus, if $U_{z_{0}}$ is a neighborhood of $z_{0}$, by [18, Lemma 11]: $P_{z_{0}}:=U_{z_{0}}\setminus W$ is pluripolar. In particular, we now show
@@
$\partial W\cap U_{z_{0}}\subseteq U_{z_{0}}\cap(\mathbb{C}^{n}\setminus W)=P_{z_{0}}.$
On the other hand, if by contradiction there exist $a\in P_{z_{0}}\cap(\mathbb{C}^{n}\setminus\overline{W})$, then there exists $U_{a}$ neighborhood of $a$ such that
$U_{a}\subset U_{z_{0}}\cap(\mathbb{C}^{n}\setminus\overline{W})\subseteq U_{z_{0}}\cap(\mathbb{C}^{n}\setminus W)=P_{z_{0}}$
which is absurd, because it is well-known (see, e.g., [4, Theorem 4.17]) that a pluripolar set has zero Lebesgue measure. It is then proved that $L_{2}$ is pluripolar, as desired. To conclude the proof, we now show $\widetilde{\Omega}\setminus E=W$. Clearly $\widetilde{\Omega}\setminus E\supseteq W$. Conversely, just notice that
$\widetilde{\Omega}\setminus E\subset W\cup(\mathbb{C}^{n}\setminus\overline{W})$
and if $(\widetilde{\Omega}\setminus E)\cap(\mathbb{C}^{n}\setminus\overline{W})\neq\emptyset$, we would have a separation of $\widetilde{\Omega}\setminus E$, which contradicts the well-known fact (see, e.g., [4, Corollary 5.26]) that a pluripolar and relatively closed subset of a domain does not disconnect the domain.